OpenStudy (anonymous):

Identify the conic section that is represented by the equation: x2 - 3y2 - 6x + 2y - 5 = 0. a. hyperbola b. Parabola c. Ellipse d. circle

OpenStudy (anonymous):

helppppp..?

OpenStudy (anonymous):

Based on that equation, you have a hyperbola. Notice that you could rearrange that to give the following: \[x^2 -3y^2-6x + 2y -5 = 0\] \[(x^2 - 6x) - 3(y^2 - \frac{2}{3}y) - 5 = 0\] \[(x-3)^2 - 3(y-\frac{1}{3})^2 - 5 -9 - \frac{1}{3} =0 \] \[(x-3)^2 - 3(y - \frac{1}{3})^2 - \frac{43}{3} = 0\] If there was a plus sign between the x and y terms, you would have an ellipse. Since it's negative, you've got an hyperbola.

OpenStudy (anonymous):

Oops, shoot, that last equation should have 41/3, not 43/3....other than that, it's fine.